The largest unsigned decimal number that can be represented in binary using six bits is
Answer #1

The largest unsigned decimal number that can be represented using 6 - bits in binary is = 111111

Here 111111 is a binary number.If we convert this into decimal form,that will be our answer.

111111 is represented in decimal as follows:

111111=(1×25)+(1×24)+(1×23)+(1×22)+(1×21)+(1×20)=63

This can also be found using the formula:

2n-1 where n is number of bits

=>26-1=64-1=63

Hence the largest unsigned decimal number that can be represented using 6-bits in binary is 63.